A Bibliography of Supercomputing ’91

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.04

Abstract

This bibliography records articles presented at the Supercomputing ’91 conference.

Title word cross-reference

15 [MS91]. 1990 [STVS91].

2600 [SWL^{+}91].

3 [SWL^{+}91].

4 [KSY91].

’91 [IEE91].

access [BC91, BS91]. accesses [GV91]. across [MMF^{+}91, NGHG^{+}91]. adaptive [FGPS91]. Albuquerque [IEE91]. algebra [KS91]. algorithm [Pug91, TG91, YG91, de 91]. algorithms [FGPS91, Nar91]. alias [Chi91]. Alleviation [FWW91]. analyses [TRM91]. Analysis [Nar91, CV91, Pug91, WL91]. animation [UHP91]. aperture [PMG91]. Application [RJCC91]. applications [DKMS91, MK91]. applied [Chr91]. approach [KS91]. approaches [TMB91]. Architecture [HQA^{+}91, AFH^{+}91, BP91a, Kar91, KSS^{+}91, Tze91]. Architecture-independent [HQA^{+}91]. architectures [BL91, Nar91, TRM91]. array [BL91, GV91, Nar91]. arrays [QM91]. arrhythmias [Sad91]. auditorialization [HW91]. automatic [CP91, KS91].

balance [VN91]. balancing [AGM91, KSY91]. bank [BU91]. based [BDG^{+}91, Kar91, KSS^{+}91]. behavior [FLK^{+}91, MK91]. Bell [DKMS91, STVS91]. benchmarks [BBB^{+}91, BF91]. benchmarks-summary [BBB^{+}91]. block [Chr91, KS91]. Boltzmann [Elt91]. bounded [Aln91]. bounded-degree [Aln91]. Burgers [Elt91].

C [HQA^{+}91, Smi91]. cardiac [Sad91]. case [HQA^{+}91]. CG [Chr91]. challenge [CFK^{+}91]. Chaotic [Sad91]. Characterizing [GPCJ91]. checkpoint [BP91b]. chip [BC91]. class [Chr91]. Climate [Che91, MMF^{+}91]. code [EOS91, HKM91]. coherence [CV91]. communicating [IAS^{+}91]. communication [VN91]. communications [Koe91, QM91]. community [MW91]. Comparison [CV91, NGHG^{+}91, SWL^{+}91]. Compiler [HKT91, Koe91, STVS91, AFH^{+}91]. Compiler-time [Koe91]. compilers [Smi91]. complex [SBGM91]. computation [Kar91]. computational [Elt91]. computations [KS91]. computer [TRM91]. computers [JH91, RJCC91]. Computing [Dan91, SBGM91, IAS^{+}91, TV91, Wil91]. concurrent [BDG^{+}91, KWFN91]. conditioned [PF91]. conflict [SN91]. conflict-free [SN91]. connected [Tze91]. Connection [MS91, SDM91]. conquer [GK91]. consistency [DWB^{+}91]. constant [Aln91]. contention [BU91]. contentions [BS91]. costs [PBH91]. course [TV91]. Cray [SWL^{+}91]. cube [Tze91]. cube-connected [Tze91]. cycles [Tze91].

DAGs [YG91]. data [BC91, GV91, Kan91, MC91, Nar91]. dataparallel [HQA^{+}91]. debate [Can91]. debugging [GH91]. decoupled [BU91]. degree [Aln91]. Delayed [DWB^{+}91]. dense [Mal91]. dependence [Pug91]. Design [Tze91, WL91, CM91, SN91]. Detecting [GV91]. detection [MC91]. development [BDG^{+}91]. differential [HIS91]. dimensional [Elt91, TRM91]. directed [BP91a]. directory [CV91]. distributed [AGM91, AFH^{+}91, HKT91, IAS^{+}91, Mal91, SY91]. distributed-memory [HKT91]. Distributing [NGHG^{+}91]. Distribution [MMF^{+}91, KSY91]. divide [GK91]. divide-and-conquer [GK91]. division [QM91, de 91]. DNA [NGHG^{+}91]. DOACROSS [CY91, SY91]. dynamic [Chi91]. dynamics [TMB91].

effective [BC91]. Effects [VN91, DWB^{+}91]. Efficient [SY91, Chr91, TG91, WL91]. element [BL91, TRM91]. elliptic [STVS91]. EM-4 [KSY91]. emerging [KSS^{+}91]. empirical [CY91, LMSS91]. environment [Dan91]. equation [Elt91, HIS91]. Euclidean [de 91]. evaluation [CP91]. examination [TG91]. execution [SY91]. experiments [LG91]. explicit [PBH91]. Exploration [Wil91].

Factoring [HSF91]. factorization

[Mal91, VN91]. fast [Pug91, YG91, de 91]. FFTs [lWR91]. finite [Kar91, TRM91]. finite-element [TRM91]. floating [MMNP91]. floating-point [MMNP91]. flow [EOS91, SBGM91]. fluid [HIS91]. fly [MC91]. fork [MC91]. fork-join [MC91]. format [MMNP91]. FORTRAN [Can91, UHP91]. FortranD [HKT91]. framework [CM91]. free [SN91]. Fujitsu [SWL^{+}91]. Fully [FGPS91]. Fully-adaptive [FGPS91]. function [KSY91]. functions [HIS91].

generation [BL91, HKM91, Koe91]. generator [STVS91]. geometries [Kar91, SBGM91]. geophysics [Wil91]. Gigaflops [MS91]. Gigascale [Mei91]. good [Smi91]. Gordon [STVS91, DKMS91]. GPFP [BL91]. Graphical [BDG^{+}91, UHP91]. grid [STVS91]. GSI [Mei91]. heterogeneous [NGHG^{+}91]. hierarchical [WL91]. High [GK91, YZ91, CFK^{+}91, CM91, MMF^{+}91]. high-performance [CM91]. high-speed [MMF^{+}91]. highly [Tze91]. hot [GPCJ91]. hypercubes [lWR91].

ill [PF91]. ill-conditioned [PF91]. image [PMG91]. implementation [Chr91, EOS91]. implications [TRM91]. independent [HQA^{+}91]. information [HW91]. Input [MK91]. Input/output [MK91]. integer [Pug91]. integrated [Chi91]. integration [Mei91]. intelligence [MW91]. Intelligent [IAS^{+}91]. interactive [CP91]. interconnection [FWW91, WL91]. Interprocedural [HKM91]. issues [RJCC91]. iteration [RS91]. iterative [PF91].

join [MC91].

K2 [AFH^{+}91].

language [HIS91]. Large [KWFN91, PBH91, RJCC91]. lattice [Elt91]. lectures [DKMS91]. level [GK91]. libraries [MMNP91]. linear [Chr91, FJ91, KS91, PF91]. Load [KSY91, AGM91, VN91]. localization [PBH91]. lookahead [BU91]. loops [CY91, HSF91]. LU [Mal91].

machine [GPCJ91, MS91, SDM91]. machines [HKT91, RS91]. management [Chi91]. mapping [IAS^{+}91]. Mass [MW91]. Massively [TV91, BL91, FLK^{+}91, RJCC91, SBGM91]. math [MMNP91]. matrix [Kar91, VN91]. Measurement [BS91]. mechanism [BP91b]. media [EOS91]. memory [AFH^{+}91, BU91, BP91b, BS91, Chi91, GPCJ91, GH91, HKT91, Mal91, PBH91, RS91, SN91, SY91, de 91]. method [Chr91, Elt91, HSF91, Kan91]. Mexico [IEE91]. mid [TV91]. mid-course [TV91]. milieu [KWFN91]. MIMD [Dan91, GPCJ91, HKT91]. miss [DWB^{+}91]. model [MMF^{+}91]. modeling [Che91, Dan91, MS91, SDM91]. molecular [TMB91]. MOVE [CM91]. MP [SWL^{+}91]. MTOOL [GH91]. multicomputer [AGM91]. multicomputers [KSS^{+}91]. multidimensional [RS91]. multiple [BS91]. multiplexed [JH91]. multiprocessor [GH91, Mal91, PBH91, QM91]. multiprocessors [SN91, SY91, ZB91]. multistage [FWW91, FJ91].

NAS [BBB^{+}91, BF91]. NEC [SWL^{+}91]. nested [MC91]. Network [KSS^{+}91, BDG^{+}91]. Network-based [KSS^{+}91, BDG^{+}91]. networks [Aln91, FWW91, FJ91, MMF^{+}91, WL91]. next [BL91]. non [CWW91]. non-uniform [CWW91]. nonshared [RS91]. November [IEE91]. number [Aln91, YG91].

Object [LG91]. Ocean [SDM91]. Omega [Pug91]. on-chip [BC91]. On-the-fly [MC91]. operand [BU91]. operating [AFH^{+}91]. optical [JH91, QM91]. Optimal [Aln91]. optimizations [HKT91]. oriented [LG91]. output [MK91].

package [PF91]. packet [FGPS91]. pairs [TG91]. Parallel [TMB91, lWR91, AFH^{+}91, BBB^{+}91, BF91, BL91, CWW91, CP91, Che91, Chr91, DWB^{+}91, EOS91, FLK^{+}91, HKM91, HSF91, Kar91, KSS^{+}91, LG91, PMG91, RJCC91, SBGM91, TG91, TV91, UHP91, Wil91]. parallelism [CWW91, GK91, MC91]. parallelization [KS91, STVS91]. partial [HIS91]. partitioning [BP91a, VN91]. patterns [Koe91]. PDEQSOL [HIS91]. penalty [BC91]. Performance [AGM91, BF91, GH91, CM91, FGPS91, SWL^{+}91, YZ91]. permutations [FJ91]. PILS [PF91]. pivoting [Mal91]. point [MMNP91]. Pointer [LMSS91]. porous [EOS91]. power [lWR91]. power-of-two [lWR91]. practical [HSF91, Pug91]. prediction [AGM91]. preliminary [BBB^{+}91]. preloading [BC91]. prime [de 91]. Prize [DKMS91, STVS91]. problem [TV91]. problems [HIS91]. Proceedings [IEE91]. processes [IAS^{+}91]. processing [BL91, Kan91, KWFN91, PMG91, YZ91]. processor [AFH^{+}91, BP91a, BS91, CM91, Nar91, SBGM91]. processors [YG91]. profiles [CWW91]. Programming [PBH91, CP91, HQA^{+}91, LG91, Pug91]. programs [CWW91, DWB^{+}91, FLK^{+}91, GH91, MC91, UHP91]. projective [Kar91]. protein [NGHG^{+}91]. prototype [KSY91].

races [MC91]. radar [PMG91]. Radix [ZB91]. range [TMB91]. rate [DWB^{+}91]. reality [Wil91]. recovery [BP91b]. reduce [BC91, BU91]. redundant [GV91]. references [BU91]. regular [Koe91]. rekindled [Can91]. reliable [Tze91]. replicated [Nar91]. requirements [MW91]. reservoir [KWFN91, RJCC91, YZ91]. resolution [Chi91]. results [BBB^{+}91, BF91, Elt91, LG91]. Retire [Can91]. robust [HSF91]. rollback [BP91b]. rounds [Aln91]. routing [FGPS91].

saturation [FWW91]. scale [KWFN91, PBH91, RJCC91]. Scheduling [CWW91, HSF91, VN91, YG91]. scheme [BC91, Chi91]. schemes [CV91]. school [CFK^{+}91]. scientific [HQA^{+}91, HW91]. Seismic [MS91]. semantics [BP91a]. semantics-directed [BP91a]. sequences [NGHG^{+}91]. shared [GPCJ91, GH91, Kan91, PBH91, SY91]. shared-memory [SY91]. short [TMB91]. simulation [KWFN91, RJCC91, YZ91]. simulations [TMB91]. software [CV91]. Solution [HIS91]. solver [HIS91, PF91]. sort [ZB91]. sorting [Aln91]. spaces [RS91]. sparse [Kar91, VN91]. speed [MMF^{+}91]. spots [GPCJ91]. static [YG91]. storage [MW91]. studies [HQA^{+}91]. study [CY91, LMSS91]. summary [BBB^{+}91]. supercomputer [BL91, DKMS91]. supercomputers [Che91, NGHG^{+}91, PMG91, SWL^{+}91]. Supercomputing [IEE91, BDG^{+}91, CFK^{+}91, MK91]. support [BP91b, GK91]. switching [FGPS91]. SX [SWL^{+}91]. SX-3 [SWL^{+}91]. symbolic [Kan91]. Synthetic [PMG91]. system [AFH^{+}91]. systems [AGM91, BS91, Chr91, IAS^{+}91, PF91, de 91].

target [LMSS91]. technology [Mei91]. test [Pug91]. Three [TRM91, HQA^{+}91, SWL^{+}91]. Three-dimensional [TRM91]. Threshold [Mal91]. Tiling [RS91]. Time [JH91, QM91, Koe91]. Time-division [QM91]. tools [BDG^{+}91, CP91]. tracking [LMSS91, TV91]. transformations [HKM91]. translation [BP91b]. tree [FWW91]. tridiagonal [Chr91]. turbulent [SBGM91]. two [BF91, Elt91, lWR91]. two-dimensional [Elt91].

ultra [de 91]. unbounded [YG91]. uniform [CWW91]. Universal [FJ91]. Using [BU91].

Vector

[EOS91, BS91, Che91, Kan91, YZ91, ZB91]. Vector/parallel [EOS91]. Vectorizing [Smi91]. via [FJ91]. virtual [BP91b]. viscous [Elt91]. Visualizing [FLK^{+}91]. VLSI [Aln91]. VP [SWL^{+}91]. VP-2600 [SWL^{+}91].

Wide [MMNP91]. wormhole [FGPS91].

Y-MP [SWL^{+}91].

References

Annaratone:1991:KDM

[AFH^{+}91] M. Annaratone, M. Fillo, M. Halbherr, R. Ruhl, Steiner, P., and M. Viredaz. The K2 distributed memory parallel processor: architecture, compiler and operating system. In IEEE [IEE91], pages 900–909. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Ahmad:1991:PPD

[AGM91] I. Ahmad, A. Ghafor, and K. Mehrotra. Performance prediction of distributed load balancing on multicomputer systems. In IEEE [IEE91], pages 830–839. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Alnuweiri:1991:OBV

[Aln91]  H. M. Alnuweiri.  Optimal

bounded-degree VLSI networks for sorting in a constant number of rounds. In IEEE [IEE91], pages 732–739. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Bailey:1991:NPB

[BBB^{+}91] D. H. Bailey, E. Barszcz, J. T. Barton, D. S. Browning, R. L. Carter, L. Dagum, R. A. Fatoohi, P. O. Frederickson, T. A. Lasinski, R. S. Schreiber, H. D. Simon, V. Venkatakrishnan, and S. K. Weeratunga. The NAS parallel benchmarks-summary and preliminary results. In IEEE [IEE91], pages 158–165. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Baer:1991:EOP

[BC91] J.-L. Baer and T.-F. Chen. An effective on-chip preloading scheme to reduce data access penalty. In IEEE [IEE91], pages 176–186. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Beguelin:1991:GDT

[BDG^{+}91] A. Beguelin, J. J. Dongarra, G. A. Geist, R. Manchek, and V. S. Sunderam. Graphical development tools for networkbased concurrent supercomputing. In IEEE [IEE91], pages 435–444. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Bailey:1991:PRT

[BF91] D. H. Bailey and P. O. Frederickson. Performance results for two of the NAS parallel benchmarks. In IEEE [IEE91], pages 166–173. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Beal:1991:GAP

[BL91] D. Beal and C. Lambrinoudakis.

GPFP: an array processing element for the next generation of massively parallel supercomputer architectures. In IEEE [IEE91], pages 348–357. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Bird:1991:SPP

[BP91a] P. L. Bird and U. F. Pleban. A semantics-directed partitioning of a processor architecture. In IEEE [IEE91], pages 702–709. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Bowen:1991:VMT

[BP91b] N. S. Bowen and D. K. Pradhan. A virtual memory translation mechanism to support checkpoint and rollback recovery. In IEEE [IEE91], pages 890–899. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Bucher:1991:MMA

[BS91] I. Y. Bucher and M. L. Simmons. Measurement of memory access contentions in multiple vector processor systems. In IEEE [IEE91], pages 806–817. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Bird:1991:ULR

[BU91] P. L. Bird and R. A. Uhlig. Using lookahead to reduce memory bank contention for decoupled operand references. In IEEE [IEE91], pages 187–196. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Cann:1991:RFD

[Can91] D. Cann. Retire FORTRAN?

A debate rekindled. In IEEE [IEE91], pages 264–272. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Cohen:1991:HSS

[CFK^{+}91] M. Cohen, M. Foster, D. Kratzer, P. Malone, and A. Solem. A high school supercomputing challenge. In IEEE [IEE91], pages 68–75. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Chervin:1991:CMP

[Che91] R. Chervin. Climate modeling with parallel vector supercomputers. In IEEE [IEE91], pages 677–?? ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Chiueh:1991:IMM

[Chi91] T.-C. Chiueh. An integrated memory management scheme for dynamic alias resolution. In IEEE [IEE91], pages 682– 691. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Chronopoulos:1991:TEP

[Chr91] A. T. Chronopoulos. Towards efficient parallel implementation of the CG method applied to a class of block tridiagonal linear systems. In IEEE [IEE91], pages 578–587. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Corporal:1991:MFH

[CM91] Corporal, H. and H. Mulder.

MOVE: a framework for highperformance processor design. In IEEE [IEE91], pages 692– 701. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Cheng:1991:EAI

[CP91] D. Y. Cheng and D. M. Pase.

An evaluation of automatic and interactive parallel programming tools. In IEEE [IEE91], pages 412–423. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Chen:1991:CAS

[CV91] Y.-C. Chen and A. V. Veidenbaum. Comparison and analysis of software and directory coherence schemes. In IEEE [IEE91], pages 818–829. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Chang:1991:SPP

[CWW91] Yao-Jen Chang, J.-L. C. Wu, and Jingshown Wu. Scheduling parallel programs with non-uniform parallelism profiles. In IEEE [IEE91], pages 502–511. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Chen:1991:ESD

[CY91] D.-K. Chen and P.-C. Yew. An empirical study on DOACROSS loops. In IEEE [IEE91], pages 620–632. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Dannevik:1991:CMM

[Dan91] W. Dannevik. Computing modeling in a MIMD environment. In IEEE [IEE91], pages 678– ?? ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

deDinechin:1991:UFE

[de 91] B. D. de Dinechin. A ultra fast Euclidean division algorithm for prime memory systems. In IEEE [IEE91], pages 56–65. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Dongarra:1991:GBP

[DKMS91] J. J. Dongarra, A. Karp, K. Miura, and H. D. Simon. Gordon Bell Prize lectures (supercomputer applications). In IEEE [IEE91], pages 328–337. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Dubois:1991:DCE

[DWB^{+}91] M. Dubois, J.-C. Wang, L. A. Barroso, K. Lee, and Y.-S. Chen. Delayed consistency and its effects on the miss rate of parallel programs. In IEEE [IEE91], pages 197–206. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Elton:1991:LBM

[Elt91] B. H. Elton. A lattice Boltzmann method for a two-dimensional viscous Burgers equation: computational results. In IEEE [IEE91], pages 242–252. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Ewing:1991:VPI

[EOS91] R. Ewing, P. O’Leary, and J. Sochacki. Vector/parallel implementation of a porous media flow code. In IEEE [IEE91], pages 294–303. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Felperin:1991:FRP

[FGPS91] S. A. Felperin, L. Gravano, G. D.

Pifarre, and J. C. L. Sanz. Fullyadaptive routing: packet switching performance and wormhole algorithms. In IEEE [IEE91], pages 654–663. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Fiduccia:1991:UMN

[FJ91] C. M. Fiduccia and E. M. Jacobson. Universal multistage networks via linear permutations. In IEEE [IEE91], pages 380–389. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Friedell:1991:VBM

[FLK^{+}91] M. Friedell, M. LaPolla, S. Kochhar, S. Sistare, Juda, and J. Visualizing the behavior of massively parallel programs. In IEEE [IEE91], pages 472–480. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Farrens:1991:ATS

[FWW91] M. Farrens, B. Wetmore, and A. Woodruff. Alleviation of tree saturation in multistage interconnection networks. In IEEE [IEE91], pages 400–409. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Goldberg:1991:PDS

[GH91] A. J. Goldberg and J. L. Hennessy. Performance debugging shared memory multiprocessor programs with MTOOL. In IEEE [IEE91], pages 481– 490. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Gursoy:1991:HLS

[GK91] A. Gursoy and L. V. Kale. High level support for divide-and-conquer parallelism. In IEEE [IEE91], pages 283–292. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Glen:1991:CMH

[GPCJ91] R. R. Glen, D. V. Pryor, J. M.

Conroy, and T. Johnson. Characterizing memory hot spots in a shared memory MIMD machine. In IEEE [IEE91], pages 554–566. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Granston:1991:DRA

[GV91] E. D. Granston and A. V. Veidenbaum. Detecting redundant accesses to array data. In IEEE [IEE91], pages 854–865. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hirayama:1991:SFP

[HIS91] H. Hirayama, M. Ikeda, and N. Sagawa. Solution functions of PDEQSOL (partial differential equation solver language) for fluid problems. In IEEE [IEE91], pages 218–227. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hall:1991:ITP

[HKM91] M. W. Hall, K. Kennedy, and K. S. McKinley. Interprocedural transformations for parallel code generation. In IEEE [IEE91], pages 424–434. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hiranandani:1991:COF

[HKT91] S. Hiranandani, K. Kennedy, and C.-W. Tseng. Compiler optimizations for FortranD on MIMD distributed-memory machines. In IEEE [IEE91], pages 86–100. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hatcher:1991:ASP

[HQA^{+}91] P. J. Hatcher, M. J. Quinn, R. J. Anderso, A. J. Lapadula, B. K. Seevers, and A. F. Bennett. Architecture-independent scientific programming in dataparallel C: three case studies. In IEEE [IEE91], pages 208–217. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hummel:1991:FPR

[HSF91] S. F. Hummel, E. Schonberg, and L. E. Flynn. Factoring: a practical and robust method for scheduling parallel loops. In IEEE [IEE91], pages 610–619. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hotchkiss:1991:ASI

[HW91] R. S. Hotchkiss and C. L.

Wampler. The auditorialization of scientific information. In IEEE [IEE91], pages 453–461. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Ieumwananonthachai:1991:IMC

[IAS^{+}91] A. Ieumwananonthachai, A. N. Aizawa, S. R. Schwartz, Wah, B. W., and J. C. Yan. Intelligent mapping of communicating processes in distributed computing systems. In IEEE [IEE91], pages 512–521. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

IEEE:1991:PSA

[IEE91] IEEE, editor. Proceedings, Supercomputing ’91: Albuquerque, New Mexico, November 18–22, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Jordan:1991:TMO

[JH91] H. F. Jordan and V. P. Heuring.

Time multiplexed optical computers. In IEEE [IEE91], pages 370–378. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Kanada:1991:MVP

[Kan91] Y. Kanada. A method of vector processing for shared symbolic data. In IEEE [IEE91], pages 722–731. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Karmarkar:1991:NPA

[Kar91] N. Karmarkar. A new parallel architecture for sparse matrix computation based on finite projective geometries. In IEEE [IEE91], pages 358–369. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Koelbel:1991:CGR

[Koe91]  C. Koelbel.  Compiler-time

generation of regular communications patterns. In IEEE [IEE91], pages 101–110. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Kung:1991:NAA

[KS91] H. T. Kung and J. Subhlok. A new approach for automatic parallelization of block linear algebra computations. In IEEE [IEE91], pages 122–129. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Kung:1991:NME

[KSS^{+}91] H. T. Kung, R. Sansom, S. Schlick, P. Steenkiste, M. Arnould, F. J. Bitz, F. Christianson, E. C. Cooper, Ö. Menzilcioğlu, D. Ombres, and B. Zill. Network-based multicomputers: an emerging parallel architecture. In IEEE [IEE91], pages 664–673. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Kodama:1991:LBF

[KSY91] Y. Kodama, S. Sakai, and Y. Yamaguchi. Load balancing by function distribution on the EM-4 prototype. In IEEE [IEE91], pages 522–531. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Kendall:1991:LSR

[KWFN91] R. P. Kendall, J. R. Wallis, J. A. Foster, and J. S. Nolen. Large scale reservoir simulation in the concurrent processing milieu. In IEEE [IEE91], pages 542–?? ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Lee:1991:OOP

[LG91] J. K. Lee and D. Gannon. Object oriented parallel programming experiments and results. In IEEE [IEE91], pages 273–282. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Loeliger:1991:PTT

[LMSS91] J. Loeliger, R. Metzger, M. Seligman, and S. Stroud. Pointer target tracking — an empirical study. In IEEE [IEE91], pages 14–23. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Woo:1991:PPF

[lWR91] M. l. Woo and R. A. Renaut. Parallel power-of-two FFTs on hypercubes. In IEEE [IEE91], pages 754–763. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Malard:1991:TPD

[Mal91] J. Malard. Threshold pivoting for dense LU factorization on distributed memory multiprocessor. In IEEE [IEE91], pages 600–607. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Mellor-Crummey:1991:ODD

[MC91] J. Mellor-Crummey. On-the-fly detection of data races for programs with nested fork-join parallelism. In IEEE [IEE91], pages 24–33. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Meindl:1991:GIG

[Mei91] J. D. Meindl. Gigascale integration (GSI) technology. In IEEE [IEE91], pages 534–538. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Miller:1991:IOB

[MK91] E. L. Miller and R. H. Katz. Input/output behavior of supercomputing applications. In IEEE [IEE91], pages 567–576. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Mechoso:1991:DCM

[MMF^{+}91] C. R. Mechoso, C.-C. Ma, J. D. Farrar, J. A. Saphr, Moore, and R. W. Distribution of a climate model across high-speed networks. In IEEE [IEE91], pages 253–260. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Markstein:1991:WFF

[MMNP91] V. Markstein, P. Markstein, T. Nguyen, and S. Poole. Wide format floating-point math libraries. In IEEE [IEE91], pages 130–138. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Myczkowski:1991:SMA

[MS91] J. Myczkowski and G. Steele.

Seismic modeling at 15 Gigaflops on the Connection Machine. In IEEE [IEE91], pages 316–326. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Myers:1991:MSR

[MW91] T. Myers and E. Williams. Mass storage requirements in the intelligence community. In IEEE [IEE91], pages 878–889. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Narayanan:1991:ARD

[Nar91] P. J. Narayanan. Analysis of replicated data algorithms on processor array architectures. In IEEE [IEE91], pages 764– 773. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM).

LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Nicholas:1991:DCD

[NGHG^{+}91] H. Nicholas, G. Giras, V. Hartonas-Garmhausen, M. Kopko, C. Maher, and A. Ropelewski. Distributing the comparison of DNA and protein sequences across heterogeneous supercomputers. In IEEE [IEE91], pages 139– 146. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Picano:1991:PCE

[PBH91] S. Picano, E. D. Brooks, III, and J. E. Hoag. Programming costs of explicit memory localization on a large scale shared memory multiprocessor. In IEEE [IEE91], pages 36–45. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Pommerell:1991:PIL

[PF91] C. Pommerell and W. Fichtner. PILS: an iterative linear solver package for ill-conditioned systems. In IEEE [IEE91], pages 588–599. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Plimpton:1991:SAR

[PMG91] S. Plimpton, G. Mastin, and D. Ghiglia. Synthetic aperture radar image processing on parallel supercomputers. In IEEE [IEE91], pages 446–452. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Pugh:1991:OTF

[Pug91] W. Pugh. The Omega test: a fast and practical integer programming algorithm for dependence analysis. In IEEE [IEE91], pages 4–13. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Qiao:1991:TOC

[QM91] C. Qiao and R. G. Melham. Time-division optical communications in multiprocessor arrays. In IEEE [IEE91], pages 644–653. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Rutledge:1991:AIL

[RJCC91] J. M. Rutledge, D. R. Jones, W. H. Chen, and E. Y. Chung. Application issues for large scale reservoir simulation on massively parallel computers. In IEEE [IEE91], pages 541–?? ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Ramanujan:1991:TMI

[RS91] J. Ramanujan and P. Sadayappan. Tiling multidimensional iteration spaces for nonshared memory machines. In IEEE [IEE91], pages 111–120. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Sadananda:1991:CCA

[Sad91] R. Sadananda. Chaotic cardiac arrhythmias. In IEEE [IEE91], pages 76–84. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Sethian:1991:CTF

[SBGM91] J. A. Sethian, J.-P. Brunet, A. Greenberg, and J. P. Mesirov. Computing turbulent flow in complex geometries on a massively parallel processor. In IEEE [IEE91], pages 230– 241. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Smith:1991:OMC

[SDM91] R. D. Smith, J. K. Dukowicz, and R. C. Malone. Ocean modeling on the Connection Machine. In IEEE [IEE91], pages 679–?? ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Smith:1991:VCC

[Smi91] L. L. Smith. Vectorizing C compilers: how good are they? In IEEE [IEE91], pages 544– 553. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Shing:1991:CMD

[SN91] H. Shing and L. M. Ni. A conflict-free memory design for multiprocessors. In IEEE [IEE91], pages 46–55. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Sabot:1991:CPE

[STVS91] G. Sabot, L. Tennies, A. Vasilevsky, and R. Shapiro. Compiler parallelization of an elliptic grid generator for 1990 Gordon Bell Prize. In IEEE [IEE91], pages 338–346. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Simmons:1991:PCT

[SWL^{+}91] M. L. Simmons, H. J. Wasserman, O. M. Lubeck, C. Eoyang, R. Mendez, H. Harada, and M. Ishiguro. A performance comparison of three supercomputers: Fujitsu VP-2600, NEC SX-3, and Cray Y-MP. In IEEE [IEE91], pages 150–157. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Su:1991:EDE

[SY91] H.-M. Su and P.-C. Yew. Efficient Doacross execution on distributed shared-memory multiprocessors. In IEEE [IEE91], pages 842–853. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Theobald:1991:EPA

[TG91] K. B. Theobald and G. R. Gao. An efficient parallel algorithm for all pairs examination. In IEEE [IEE91], pages 742–753. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Tamayo:1991:PAS

[TMB91] P. Tamayo, J. P. Mesirov, and B. M. Boghosian. Parallel approaches to short range molecular dynamics simulations. In IEEE [IEE91], pages 462– 470. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Taylor:1991:TFA

[TRM91] V. E. Taylor, A. Ranade, and D. G. Messerschmitt. Threedimensional finite-element analyses: implications for computer architectures. In IEEE [IEE91], pages 786–795. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Tomkins:1991:MPC

[TV91] J. L. Tomkins and J. P.

VanDyke. Massively parallel computing and the mid-course tracking problem. In IEEE [IEE91], pages 796–804. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Tzeng:1991:DHR

[Tze91]  N.-F. Tzeng.  Design of a

highly reliable cube-connected cycles architecture. In IEEE [IEE91], pages 776–785. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Utter-Honig:1991:GAP

[UHP91] S. Utter-Honig and C. M. Pancake. Graphical animation of parallel Fortran programs. In IEEE [IEE91], pages 491– 500. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Venugopal:1991:EPS

[VN91] S. Venugopal and V. K. Naik. Effects of partitioning and scheduling sparse matrix factorization on communication and load balance. In IEEE [IEE91], pages 866–875. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Willen:1991:EGP

[Wil91] D. E. Willen. Exploration geophysics, parallel computing and reality. In IEEE [IEE91], pages 540–?? ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Wei:1991:DAE

[WL91] S. Wei and S. Levy. Design and analysis of efficient hierarchical interconnection networks. In IEEE [IEE91], pages 390–399. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Yang:1991:FSS

[YG91] T. Yang and A. Gerasoulis. A fast static scheduling algorithm for DAGs on an unbounded number of processors. In IEEE [IEE91], pages 633–642. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Young:1991:HPV

[YZ91] L. C. Young and S. E. Zaratonello. High performance vector processing in reservoir simulation. In IEEE [IEE91], pages 304–315. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Zagha:1991:RSV

[ZB91] M. Zagha and G. E. Blelloch.

Radix sort for vector multiprocessors. In IEEE [IEE91], pages 712–721. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.